the sum of the ages of mike and dan is 20 years. in 4 years, dan's age will be three-fourths of mike's age. how old is mike now?

Answer by Anthea Lawn(22)

let m = mikes age, d = dans age.
from the first statement : m+d = 20 so d = 20-m
from the second statement (3/4)(m+4) = d+4
working a bit on the second statement gives :
3(m+4) = 4(d+4)
3m + 12 = 4d + 16
3m = 4d + 4
substituting from the first statement :
3m = 4(20-m) + 4
3m = 80 - 4m + 4
7m = 84
m = 12
So Mike is 12 and Dan is 8